Nationwide, 65 percent of persons taking a certain professional certification exam pass. Consider, for a samples of 200, the sampling distribution of ?̂ p^. (Each answer should be entered as a proportion rounded to three decimal places.)

The standard deviation of the sampling distribution of ?̂ p^ is:
Answer #1

Solution

Given that,

p = 65%=0.65

1 - p = 1-0.65=0.35

n = 200

= p =0.65

The standard deviation of the sampling distribution of ?̂ p^ is

=  [p ( 1 - p ) / n] = [(0.65*0.35) / 200 ] = 0.034